OpenStudy (anonymous):

f(x) = 6/7x ; f '(x) =?

OpenStudy (anonymous):

the derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\) the \(\frac{6}{7}\) is a constant, leave it there